# On Robustness of Vision-Language-Action Model against Multi-Modal Perturbations

### v3

@@ 5.3 Real-World Experiments
Real-world setup. We deploy VLAs on a Fairino FR5 robotic arm with a Hitbot Z-EFG-C50 gripper. Visual inputs are provided by a ZED2 external camera and an Intel RealSense 435i wrist-mounted camera. We design four tasks to assess real-world performance: (1) pick up the blue bowl on the blue plate; (2) pick up the bowl on the plate with randomized colors; (3) place the bread on the plate; and (4) place the green cup next to the green plate. These tasks test VLA’s capability on basic grasping, semantic generalization, deformable object manipulation, and spatial reasoning. For each task, we fine-tune $\pi_{0}$ with 25 demonstration trajectories.
@@
Real-world performance. As shown in Fig. 5, RobustVLA is highly effective in real-world deployment, surpassing the best baseline by 65.6% in success rate. We attribute this to the limited number of real-world trajectories available for fine-tuning. Unlike large and diverse simulation datasets such as LIBERO, laboratory-collected data are costly and offer limited state coverage, leading baseline VLAs to overfit demonstrations and fail under perturbations. In contrast, by anticipating noise during training, RobustVLA anticipates potential noise during training and remains robust to unexpected real-world disturbances. See per-task results in Appendix C.2.
Failure Analysis. Finally, we analyze the failure mode of baselines, aggregating $\pi_{0}$ and BYOVLA due to their similar behaviors. As illustrated in Fig. 5: (1) under action uncertainties, baselines exhibit imprecise control, resulting in the gripper unable to grasp the bowl precisely and occasionally knocking the bowl over; (2) under observation uncertainties, input noise obscures the visual signal, making the robot unable to identify the object clearly; (3) under environment shifts, the added neon light produces a severe illumination change that renders observations OOD; and (4) under instruction noise, speech parsing errors lead to misinterpretation, making robots placing the cup “beside” the table. In contrast, RobustVLA remains reliable in most trials, achieving consistent success in the real world. Demo videos available in our codebase.
@@ 6 Conclusion
In this paper, we evaluate and enhance the robustness of VLA models under multi-modal uncertainties. We evalute mainstream VLAs against 17 perturbations across four modalities, showing that actions are the most fragile, existing visual-robust VLAs fail to generalize beyond vision, and $\pi_{0}$ the most robust backbone with a diffusion-based action head. Building on these insights, we propose RobustVLA, a unified framework against input and output perturbations. For output robustness, we perform offline optimization against worst-case action noise that maximizes flow mismatch. For input robustness, we enforce consistent actions across semantically equivalent inputs and automatically identify the most harmful perturbation using UCB. On LIBERO, RobustVLA delivers absolute gains of 12.6% on $\pi_{0}$ and 10.4% on OpenVLA across 17 perturbations, achieves 50.6x faster inference than visual-robust VLAs, and improves robustness by 10.4% under mixed perturbations. In real-world tasks with limited demonstrations, RobustVLA achieves absolute gains of 65.6% under four modalities of perturbations.
@@ 8 Reproducibility statement
We have open sourced all our code for robustness evaluation and enhancement in https://anonymous.4open.science/r/RobustVLA-283D. Additional experiment details are available in Appendix A and B.
@@ A.2 Observation Uncertainties
Table 3: Summarization of Observation Uncertainties
@@ C.1 Details of OpenVLA Experimental Results
C.2 Details of Real-World Experimental Results
@@
Action Uncertainty. The baseline often suffers from imprecise control, such as jitter and control errors, leading to misalignment in grasping and frequent task failures. In contrast, RobustVLA exhibits strong tolerance to such disturbances, enabling stable task execution. As shown in Fig. 5, while the baseline exhibits large-amplitude oscillations and severe localization deviations, RobustVLA reduces jitter to a minor level and consistently completes the tasks.
Observation Uncertainty. The baseline frequently misjudges task states under obscure inputs. For example, due to visual interference, it often fails to recognize task completion and repeatedly attempts to grasp the target. Although BYOVLA helped reduce distraction from irrelevant objects or background clutter, it remains insufficient for reliable state estimation. RobustVLA, however, accurately infers execution states under degraded observations, which is critical for task success.
Environment Uncertainty. The baseline often encountered OOD situations under environmental changes such as distractor objects and varying illumination, leading to degraded tracking of targets, misjudgment of self-states, or even catastrophic failures (e.g., descending directly toward the table). Our approach maintains strong target-tracking and state-estimation capabilities under diverse environmental factors.
Instruction uncertainty. Variations in object references, actions, or prepositions—through synonym substitution, word-order changes, or irrelevant content—often cause the baseline to misinterpret instructions, leading to failures in object tracking, spatial reasoning, and instruction following. RobustVLA, in contrast, consistently extracts key information from diverse instructions and executes them reliably.

### v4

@@ 5.3 Real-World Experiments
Real-world setup. We deploy VLAs on a Fairino FR5 robotic arm with a Hitbot Z-EFG-C50 gripper. Visual inputs are provided by a ZED2 external camera and an Intel RealSense 435i wrist-mounted camera. We design four tasks to assess real-world performance: (1) pick up the blue bowl on the blue plate; (2) pick up the bowl on the plate with randomized colors; (3) place the bread on the plate; and (4) place the green cup next to the green plate. These tasks test VLA’s capability on basic grasping, semantic generalization, deformable object manipulation, and spatial reasoning. For each task, we fine-tune $\pi_{0}$ with 25 demonstration trajectories. All methods achieve 100% success without perturbation after fine-tuning.
@@
Real-world performance. As shown in Fig. 5, RobustVLA is highly effective in real-world deployment, surpassing the best baseline by 65.6% in success rate. We attribute this to the limited 25 demonstrations of real-world trajectories available for fine-tuning. Unlike large and diverse simulation datasets such as LIBERO, laboratory-collected data are costly and offer limited state coverage, leading baseline VLAs to overfit demonstrations and fail under perturbations. In contrast, by anticipating noise during training, RobustVLA anticipates potential noise during training and remains robust to unexpected real-world disturbances. See per-task results in Appendix C.6.
Figure 6: Results with more demonstrations. Although $\pi_{0}$ becomes more robust as additional data is provided, RobustVLA already achieves strong robustness in the low-data regime and consistently outperforms $\pi_{0}$ across all data scales.
Failure Analysis. Finally, we analyze the failure mode of baselines, aggregating $\pi_{0}$ and BYOVLA due to their similar behaviors. As illustrated in Fig. 5: (1) under action uncertainties, baselines exhibit imprecise control, resulting in the gripper unable to grasp the bowl precisely and occasionally knocking the bowl over; (2) under observation uncertainties, input noise obscures the visual signal, making the robot unable to identify the object clearly; (3) under environment shifts, the added neon light produces a severe illumination change that renders observations OOD; and (4) under instruction noise, speech parsing errors lead to misinterpretation, making robots placing the cup “beside” the table. In contrast, RobustVLA remains reliable in most trials, achieving consistent success in the real world. See demo videos in our codebase.
Results with More Demonstrations. We evaluate our performance with 50 and 100 demonstrations in Task 1 to show results with sufficient data. While all methods achieves 100% success rate, as shown in Fig. 6, $\pi_{0}$ gains robustness with additional data. The success rate of $\pi_{0}$ increases from 37.5% at 25 demos to 60.0% at 50 demos and 65.0% at 100 demos. This shows real-world robustness benefits from additional trajectories with higher data diversity. Comparing with baselines, our RobustVLA is highly robust in low-data regime, and consistently outperforms $\pi_{0}$ across all data scales. Our success rate reaches 92.5% at 25 demos, 92.5% at 50 demos and 95.0% at 100 demos, while the baseline plateaus at 65.0%, achieving 30% robustness gain than $\pi_{0}$, showing RobustVLA provides substantial robustness gains beyond what is achievable through additional demonstrations alone.
@@ 6 Conclusion
In this paper, we evaluate and enhance the robustness of VLA models under multi-modal uncertainties. We evalute mainstream VLAs against 17 perturbations across four modalities, showing that actions are the most fragile, existing visual-robust VLAs fail to generalize beyond vision, and $\pi_{0}$ the most robust backbone. Building on these insights, we propose RobustVLA, a unified framework against input and output perturbations. For output robustness, we perform offline optimization against worst-case action noise that maximizes flow mismatch. For input robustness, we enforce consistent actions across semantically equivalent inputs and automatically identify the most harmful perturbation using UCB. On LIBERO, RobustVLA delivers absolute gains of 12.6% on $\pi_{0}$ and 10.4% on OpenVLA across 17 perturbations, achieves 50.6x faster inference than visual-robust BYOVLA, and improves robustness by 10.4% under mixed perturbations. Real-world FR5 robot under perturbation shows our RobustVLA is highly robust with limited data, outperforming $\pi_{0}$ by $65.6\%$ success rate with 25 demos. With sufficient demos, our RobustVLA continues to maintain high robustness, achieving 30% higher success rate than $\pi_{0}$.
7 Acknowledgement
This study was supported in part by the National Natural Science Foundation of China Project No. 62476018 and 62322318 and in part by the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government via the Hong Kong Centre for Logistics Robotics, and in part by National Key R&D Program of China Project 2022ZD0161100.
@@ 9 Reproducibility statement
We have open sourced all our code for robustness evaluation and enhancement in https://github.com/gakakulicc/RobustVLA. Additional experiment details are available in Appendix A and B.
@@ A.2 Observation Uncertainties
Table 3: Summarization of Observation Uncertainties.
@@ B.1.3 Adversarial Training Parameters
B.2 Discussion on Parameter Selection
Table 8: Average success rate (%) of the ablation study on the PGD action space $\epsilon$ parameter.
$\pi_{0}$ $\epsilon=0.015$ $\epsilon=0.03$(used) $\epsilon=0.06$
Clean 96.0 95.8 95.5 93.9
Uniform Noise 63.5 66.0 69.8 68.1
Gaussian Noise 31.4 37.1 36.0 35.9
Action Bias 23.0 38.5 42.3 43.7
Random Flips 52.7 56.5 58.7 56.8
Sudden Spikes 51.7 61.1 59.7 61.5
Average 44.5 51.8 53.3 53.2
Table 9: Average success rate (%) of the ablation study on the UCB parameters.
Ours exp_coef=1.5 ema=0.6 win=200
Clean 95.5 93.8 95.5 96.2
Gaussian Noise 93.8 93.9 93.6 93.7
Dead Pixel 93.8 94.0 93.4 93.6
Motion Blur 95.5 95.6 95.2 95.3
Color Jitter 69.5 72.2 68.0 70.7
Image Rotation 94.4 95.0 93.8 94.6
Image Shift 92.7 93.5 92.0 93.0
Average 90.0 90.7 89.3 90.2
Our method involves two groups of hyperparameters: (1) the PGD-based adversarial action perturbation parameters, and (2) the UCB-based adaptive noise selection parameters. As stated in Appendix. B.1, the overall performance of RobustVLA is not highly sensitive to these hyperparameters. Nevertheless, we provide a more detailed discussion on initialization and tuning principles.
B.2.1 PGD Parameters
The parameters pgd_alpha and pgd_steps directly control the optimization strength of adversarial action search. In general, larger pgd_steps allow the perturbation to approach a stronger worst-case solution. However, this improvement comes with increased computational cost, and we recommend practitioners balance effectiveness and resource constraints—using as many steps as feasible within their budget.
The perturbation magnitude adv_epsilon is less intuitive to reason about, and its sensitivity requires empirical validation. Table 8 compares multiple values of $\epsilon$ and shows that the success rates remain stable across a broad range of values; all choices outperform the non-adversarial baseline by a significant margin. This confirms that adv_epsilon is not sensitive in practice. A value around 0.03 consistently provides strong robustness while avoiding over-perturbation, and we use it as the default.
Before introducing tuning guidelines, we provision the additional ablation results in Table 9. The UCB module shows strong robustness across different hyperparameter choices: varying the ucb_exploration_coeff, ucb_ema_decay, or ucb_window_size only leads to minor fluctuations in success rate. Larger ucb_exploration_coeff slightly improve performance on harder corruptions, while smaller ucb_ema_decay values make the selection more reactive but somewhat less stable. Enlarging the ucb_min_samples provides mild stability gains. These results confirm that our default configuration offers a well-balanced trade-off without requiring precise tuning.
For the UCB-based adaptive perturbation selection, a good starting point is the configuration reported in Appendix. B.1, which has demonstrated stable performance across diverse tasks and typical noise levels.
During tuning, we recommend a data-driven, stability-first strategy. If one observes high-frequency oscillation in the training loss or overly frequent switching in selected perturbations, it is advisable to increase stability by lowering ucb_exploration_coeff, lowering ucb_ema_decay, or enlarging ucb_window_size. Conversely, if the loss plateaus for an extended period or the selected perturbations become overly concentrated—indicating insufficient adaptation—these parameters can be adjusted in the opposite direction to encourage exploration. This closed-loop adjustment ensures a balance between adaptability and convergence stability.
B.3 Training loss
Figures.  7 and 8 report the training loss curves of $\pi_{0}$ under our hyperparameter configuration, using both LIBERO simulation data and real-world trajectories. Across all loss components—including the clean loss, input-consistency loss, and adversarial output loss—the curves remain smooth and stable throughout training. This indicates that the chosen PGD parameters and UCB-based perturbation scheduling provide a reliable optimization landscape without inducing instability or oscillatory behavior.
Given this stability, we consider the presented hyperparameter setting to be a robust default for training RobustVLA. If further tuning is required for new tasks or environments, practitioners can use these loss patterns as a reference:
Excessive fluctuations or abrupt jumps in any loss component typically suggest the need to reduce the UCB exploration coefficient, increase the window size, or lower the PGD step size.
Prolonged plateaus may indicate insufficient exploration or under-strength adversarial perturbation, in which case moderate increases to these parameters can help.
Figure 7: Loss during training with LIBERO data
Figure 8: Loss during training with real-world data
@@ C.1 Details of OpenVLA Experimental Results
C.2 Details and discussion on robustness testing of $\pi_{0}$ and $\pi_{0}$-FAST
In addition to the Fig. 3(c) presented in the main paper, we further compare $\pi_{0}$ and $\pi_{0}$-FAST under progressively increasing perturbation strengths. Although these two models share the same pre-trained VLM backbone and a similar policy architecture, they differ fundamentally in their action-generation mechanisms: $\pi_{0}$ employs a diffusion-based action expert, whereas $\pi_{0}$-FAST produces actions autoregressively via token prediction.
Figure 9: $\pi_{0}$ and $\pi_{0}$-FAST under varying noise level
Figure 10: $\pi_{0}$ vs RobustVLA in LIBERO-long
To examine whether this distinction leads to different robustness characteristics, we conduct stress tests on the two most fragile modalities identified in Fig. 3(a): action perturbations and observation perturbations. For each modality, we gradually increase the perturbation magnitude and record the resulting drop in success rate.
As shown in Fig. 10, the diffusion-based $\pi_{0}$ consistently outperforms the autoregressive $\pi_{0}$-FAST across nearly all noise levels. The performance gap widens as perturbations become more severe, reinforcing our main-paper conclusion regarding robustness differences across VLA model classes. These results provide additional evidence that diffusion-based policies possess stronger intrinsic robustness to multi-modal disturbances than their autoregressive counterparts.
C.3 Details of long-horizon task Results
When evaluating complex and long-horizon tasks, LIBERO-long in the LIBERO suite as an appropriate benchmark. Therefore, we conduct an additional analysis in this section. As shown in Fig. 10, we observe that the inherently fragile action modality of the VLA $\pi_{0}$ deteriorates further on long-horizon tasks, while the instruction modality experiences a substantial drop due to compounding instruction-following errors. The robustness of the observation and environment modalities also declines to varying degrees. In contrast, RobustVLA maintains strong gains in the observation and environment modalities (39.93% and 12% on average, respectively), and further improves robustness in the more vulnerable action and instruction modalities (8% and 4.33% on average, respectively). Overall, RobustVLA achieves an average improvement of 19.61% across all four modalities on LIBERO-long.
Figure 11: $\pi_{0}$ and RobustVLA under varying noise level.
Figure 12: Effect of Flow Matching Loss on robustness of action modality.
C.4 Performance of RobustVLA under varying noise levels.
In the main text, we reported robustness improvements across the full set of uncertainty types and modalities. In this appendix subsection, we further examine how RobustVLA behaves within a single modality under varying noise intensities. We focus on the two most fragile modalities identified in our evaluation, action and observation, and systematically sweep their corresponding noise levels. As shown in Figure 12, RobustVLA consistently achieves positive robustness gains across all tested noise levels, with average improvements of 5.6% in the action modality and 23% in the observation modality. Moreover, we observe a clear trend in which robustness gains increase with noise intensity. For example, in the action modality, the improvement is only 1.3% at a noise level of 0.5% but rises to 8.7% at 5%. A similar pattern appears in observation perturbations, where the gain grows from 1.9% at a noise level of 3.9% to 74.7% at 35%. These results suggest that RobustVLA not only stabilizes performance under mild disturbances but also provides increasing benefits as perturbations become more severe.
C.5 Correlation Between Adversarial Flow-Matching Loss and Action Robustness
To assess whether flow-matching loss serves as a meaningful proxy for action robustness, we analyze the correlation between our adversarial flow-matching loss and task success under action perturbations. Conceptually, flow-matching provides a differentiable surrogate for action quality, consistent with prior work in adversarial robustness where differentiable losses are used as proxies for task-level objectives that are not directly differentiable.
In VLA imitation-learning settings, credit assignment is particularly challenging due to high-frequency control. As a result, flow-matching naturally quantifies the deviation between the model’s predicted actions and expert demonstrations. Larger flow-matching deviations correspond to larger mismatches from the demonstrated trajectories, which intuitively reduce the likelihood of successful task execution.
Following this motivation, we measure the adversarial flow-matching loss defined in Eq. 4 and compare it with the corresponding success rates under action perturbations. As shown in Fig. 12, the two quantities exhibit a clear monotonic trend. Quantitatively, the Pearson correlation coefficient is r=0.953r=-0.953 with $p<0.05$, indicating strong and statistically significant negative correlation. Intuitively, as adversarial flow-matching loss decreases (0.0686 $\rightarrow$ 0.0240), the success rate increases correspondingly (44.5% $\rightarrow$ 53.3%).
These results confirm that adversarial flow-matching loss is a reliable and interpretable indicator of action robustness: larger worst-case deviations in the action space directly translate into lower task success rates.
C.6 Details of Real-World Experimental Results
@@
Action Uncertainty. The baseline models exhibit pronounced sensitivity to action noise, where even minor deviations—such as motor calibration errors, communication jitters, or actuator delays—cause large downstream failures. In practice, once a single action deviates from the intended trajectory, the autoregressive or imitation-trained policy is unable to recover, as the SFT data seldom contain recovery demonstrations or state-reset behaviors. This results in misaligned grasps, gripper jitter, or cumulative drift of end-effector poses. As shown in Fig. 5, the baseline often displays large-amplitude oscillations and severe localization offsets.
In contrast, RobustVLA explicitly optimizes output robustness (Eq. 4) by injecting worst-case action perturbations into the flow-matching objective during training. This exposes the model to both clean and noisy action distributions, effectively smoothing the action manifold and preventing brittle, single-point failures. As a result, RobustVLA produces significantly more stable controls and suppresses error accumulation from individual noisy actions.
Observation Uncertainty. Under visual corruptions such as Gaussian noise, dead pixels, motion blur, or illumination changes, baseline models frequently misinterpret object boundaries or execution states. For example, in tasks like “place the bread on the plate,” motion blur can obscure the bread–plate contact, causing the baseline to repeatedly attempt a grasp even after the object has been placed. Although BYOVLA reduces distraction from irrelevant regions, its robustness is largely confined to static visual artifacts and remains insufficient for temporal or semantic consistency.
RobustVLA mitigates these issues through input-robustness regularization (Eq. 9), which enforces consistent actions under semantically preserving perturbations. Combined with our UCB-based adaptive perturbation scheduling, the model is trained against the most harmful visual corruptions, enabling effective feature denoising and context-aware reasoning. Consequently, RobustVLA reliably infers state transitions and target positions even under severely degraded observations.
Environment Uncertainty. Environmental variations such as distractor objects, background changes, or dynamic illumination induce OOD shifts that significantly degrade baseline performance. Because imitation-based policies tend to overfit the limited demonstration distributions, they often mis-track objects, attend to irrelevant items, or misjudge their own pose when confronted with unseen layouts. In real-world tests, such failures occasionally culminate in catastrophic behaviors (e.g., descending directly toward the table when shadows alter perceived depth).
Our framework alleviates this failure mode by prioritizing perturbations that maximally increase the flow-matching loss, ensuring that the model repeatedly encounters difficult OOD scenarios during training. Exposure to these diverse perturbations improves target tracking and state estimation under environmental variations, resulting in robust execution even when scene layouts deviate significantly from training conditions.
Instruction Uncertainty. Baseline models are highly sensitive to linguistic variations such as synonym substitutions, word-order changes, or extraneous modifiers. Since these models often rely on shallow keyword matching rather than semantic grounding, injected linguistic noise frequently alters their interpretation of spatial relations or object references. For instance, substituting “put the cup next to the plate” with “gently place the cup by the side of the dish” causes the baseline to misidentify the target position or confuse object roles, leading to failed execution.
RobustVLA addresses this issue by enforcing cross-perturbation action consistency for language inputs as well. By treating semantically equivalent linguistic variants as invariances during training, RobustVLA learns to extract stable task-relevant cues rather than brittle lexical patterns. This significantly improves instruction-following reliability under natural linguistic variability commonly observed in real human–robot interaction.